OpenStudy (mboo):

The multiplicative inverse of -3 is _____.

OpenStudy (joshtray):

\[-3\times \frac{ 1 }{ -3 } = 1\] The multiplicative inverse of a number multiplied by that number equals 1. So the answer is 1/-3

OpenStudy (mboo):

Thank You!:)

OpenStudy (joshtray):

your welcome

OpenStudy (joshtray):

Do you mean ascending or descending order? What is ()2? Don't really get this question.

OpenStudy (mboo):

I copied the question wrong. It is on online learning (Odyssey-ware) Which of the following shows the correct order for -1, (-1)2, -1.4 and ()2? A. (-1)^2 < -1.4 < -1 < (1/2)^2 B. -1.4 < -1 < (1/2)^2 < (-1)^2 C. -1 < -1.4 < (1/2)^2 < (-1)^2

OpenStudy (mboo):

(1/2)^2

OpenStudy (joshtray):

the answer is B because (1/2)^2=1/4 and (-1)^2=1

OpenStudy (mboo):

Thank you so so much!

OpenStudy (joshtray):

You'll get\[-1.4<-1<1/4<1\]